THE ELECTRONIC SPECTRUM OF CoCl_{2} IN THE GAS PHASE

\underline{P. J. HODGES}, J. M. BROWN, Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ, UK; T. D. VARBERG, Macalester College, Chemistry Department, 1600 Grand Avenue, St. Paul, MN 55105-1899.

The 310 nm uv band system of the CoCl_{2} radical has been recorded at high resolution by laser excitation spectroscopy. (The experimental details are described in full elsewhere ^{a}). The molecule was formed in the high temperature reaction between HCl(g) and cobalt metal and was cooled to a rotational temperature of about 10K in a subsequent free-jet expansion. This work confirms earlier observations by DeKock and Gruen^{b} and provides much more structural information. Excited state vibrational progressions of approximately 200 cm^{−1} have been identified and tentatively assigned to the symmetric stretching vibration. Rotational analyses of the three longest wavelength bands reveal that the lower and upper electronic states both have Ω = 7/2; this is consistent with recent ab initio calculations which predict a ^{4}Δ ground state c. The following rotational constants were determined for the (200) band: B′′ = 0.056432(81) cm^{−1}, B′ = 0.04951(10) cm^{−1}. A study by dispersed fluorescence shows progressions in the ground state symmetric stretching vibration (360 cm^{−1}). This is very similar to the value determined for other transition metal dichlorides ^{d}. Note that the suggested value for ν_{1} in the excited state is much smaller than this.

aS. H. Ashworth, F. J. Frieman, and J. M. Brown, J. Chem. Phys., 104, 48, (1996).

bC. W. DeKock and D. M. Gruen, J. Chem. Phys., 44, 4387, (1966).

cV. V. Sliznev, N. Vogt, and J. Vogt, The Eighteenth Colloquium on High Resolution Molecular Spectroscopy, Dijon , Poster O4, (2003).

dF. J. Grieman, S. H. Ashworth, and J. M. Brown, J. Chem. Phys., 92, 6365, (1990).